OpenStudy (anonymous):

write the log equation in exponential form" In 4= 1.3862 In e^3= 3 and In 1/e^2 = -2 Help Please?

OpenStudy (anonymous):

just to help you out, ln is basically log (base e)...therefore for the first problem its like we've the other problems we've been through, its e ^1.3862 = 4

OpenStudy (anonymous):

now the second one is interesting. because we have e ^3 = e ^3, the answer is simply the exponent because the bases are the same, therefore 3.

OpenStudy (anonymous):

second one is the same story, because its 1/e^2, simply change it to e^-2 since the negative sign means that the value is stored as a denominator., and the answer becomes -2....so basically just write out e^-2 = e^-2

OpenStudy (anonymous):

hey js chat with us

OpenStudy (anonymous):

does that help amy?

OpenStudy (anonymous):

Yess very much! im starting to understand all of it!

OpenStudy (anonymous):

good :) glad to hear!, this kind of stuff i rllly sucked at at first, but u gotta practice it a lot and eventually u get the hand of it...its my first year in uni (well now over) and sitll i had trouble with it and i started this pellet in 10th grade.

OpenStudy (anonymous):

hahaha awesome! haha so this will take me years haha

OpenStudy (anonymous):

not saying that! haha just saying its hard so dont kill urself over not understanding it!

OpenStudy (anonymous):

haha gotcha!

OpenStudy (anonymous):

anyway if u have more questions amy ill be back in 20 min. gl with ur studies.

OpenStudy (anonymous):

thanks!